OpenStudy (anonymous):

What is the slope, and what is the y-intercept of the line? y=1/2x-5 Slope= 1/2 y-intercept = ????

OpenStudy (anonymous):

@tburn25

OpenStudy (photon336):

Lines are in the form y = m(x)+b where m is the slope. in our line we have m = (1/2) right? but to find the y intercept we need to set x = 0, or rather plug in x = 0 into our equation. let's do that. y = 0.5(x)-5 y = 0.5(0)-5 = 0-5 y = -5 so our y intercept would be -5, why? because this is the point where our function touches the y axis at -5 when x is zero. sometimes the b value gives us the y intercept.

OpenStudy (anonymous):

The equation of any straight line, called a linear equation, can be written as: y = mx + b, where m is the slope of the line and b is the y-intercept. The y-intercept of this line is the value of y at the point where the line crosses the y axis.